Jamie Walker  jamie@sagaxis.co.uk

1 Yew House, 2 Woodland Crescent, London. SE16 6YH Mobile: 07736 781 390

Current Status

I have six years commerical experience working in capital markets more recently specializing in front office trading systems. I have extremely strong Java skills and am routinely involved in architectural decisions, however my technical skillset is broad including Perl, SQL and C.

Experience November 2007 – Present UBS, Equities Algorithmic Trading, London Associate Director

As a senior Java developer on the team, I have worked on all aspects of UBS’s market leading European algorithmic trading platform including macro and micro strategies and also smart order routing. Working closely with the quant desk and traders, I have implemented new single order and portfolio based strategies and delivered enhancements to existing functionality and client customizations involving front-to-back changes.

The platform is multi-threaded built using core Java and makes extensive use of NIO to communicate to UBS’s proprietary order management and market data systems. As part of my role, I have made core changes to the framework to improve performance, reduce latency and remove legacy dependencies.

I have also worked on UBS’s North American platform, aiding in porting strategies from legacy C++ platform onto a new strategic Java framework.

As part of a strategy to produce a unified global product, I am intimately involved in specifying the future architecture for the system and currently building an initial prototype for the next generation platform to replace encumbant systems in Europe and North America. This has led me to gain some understanding of the differeces in market structure between regions and given insight into two entirely different approaches to algo trading.

I have had direct management responsibility for junior team members and regularly coached new members of the team. I also spent some time in the US introducing the developers and quant desk to the concepts behind our trading models in Europe.

July 2004 – November 2007 Detica (formerly Evolution Consulting), London Senior Consultant

During my time at Detica, I worked as a lone consultant within client teams and delivered solutions developed in-house. I progressed rapidly from a junior developer to more senior roles, becoming more involved with architectural design and requirements specification. I delivered numerous technical presentations internally and took on responsibility for organising a monthly technical forum. I was also involved in the recruitment process: interviewing permanent and non-permanent hires and performing due diligence on candidates before an offer was confirmed.

• Liquidity Hub, Interest Rate Swaps Trading Platform: Took on a key role within a team of thirty, working closely with the lead technical architect to define and implement the middleware, transaction and configuration framework for an interest rate swaps trading platform commissioned by a consortium of more than 15 tier-one investment banks. Responsible for the middleware. Subject matter expert on JMS and Spring which formed the basis for the platform. Integration manager: defined and implemented the configuration and framework to enable the individual components to work together. This also involved co-ordinating the three main development streams to ensure coherence. Also respsonsible for tuning and testing of all non-functional performance and latency requirements.

• Goldman Sachs, Securities Lending: Worked closely with the operations technology and business teams. Delivered new functionality and small fixes to a very large securities lending trading platform making extensive use of J2EE, Sybase, JMS messaging, Autosys and Unix scripting. Responsibilities included adding new functionality to core client and server side Java components, responding to production issues, managing the StockLoan side of UATs being conducted at the firm-wide level, gathering requirements to replace legacy C/C++ X-Windows applications, performance tuning of Sybase queries, triggers and stored procedures, and coaching new members of the team. My knowledge of the system enabled me to participate in the re-design of the core trade model, for which I designed and delivered a real-time translation layer to publish trades and trade events from the old system to its replacement.

• Lehman Brothers, Equity Strategy: Designed and built a web portal to showcase the extensive database built by the Strategy Group. This included interactive charts generation and integration with the banks legacy document management system. The system was built using Java with Spring on a WebLogic 8.1 cluster.

• Lehman Brothers, Equity Research: Worked on productionizing a global research authoring and workflow solution (VB/VBA/XSLT) for Equity Research in a large international investment bank. Responsible for producing releases and implementing new features to bring the global system in line with existing regional systems.

• UBS, FX: Designed a benchmark to prove scalability of Caplin for streaming prices during prototyping of strategic FX trading platform.

Also designed reliable trade-submission protocol on top of the Caplin streaming model.

• Royal Bank of Canada, FX: Designed and built additional functionality for additional benchmarking of FX trades on client-facing (trustees) web-application (Weblogic/JSP) for a large international investment bank. Responsible for integration of new data feed and the testing, release and deployment.

• Royal Bank of Canada, FX: Made additions to frontend FX trade aggregating application (Weblogic/Swing) for large international investment bank. Responsible for testing, the release process and deployment.

June 1999 – April 2004 Clinical Audit, West Cumberland Hospital Systems Administrator / Software Developer

• Developed Clinical database system initially in Microsoft Access, and later with PostgreSQL on GNU/Linux (client-server architecture)

• Developed a web-based front-end to the clinical database in mod perl, including a Java applet to browse the Read Code thesaurus, and a Unix daemon to provide data to the Java applet

• Integrated the clinical database system with legacy systems in the hospital (e.g. to download demographic details)

• Developed a solution using Teleforms for capture of information

• Taught colleagues the ‘Perl’ programming language

• Administered Debian GNU/Linux servers and Novell Netware (3.12) servers, including migration from the latter to the former

• Installation of the Novell Netware server, including the first implementation of email within the hospital

January 2003 – December 2003 Magdalene College JCR, Cambridge Academic Affairs & IT Officer

• Designed and built an interactive website for the student body

• Responsible for the general IT needs of the committee

IT Skills

• Programming Languages: Java, Perl, VBA, VB, SQL, XML, XSLT, HTML, Javascript, CSS, C/C++ (both on Win32 and POSIX), L^{A}TEX

• Technologies: WebLogic (inc. JRockit Realtime), Spring, FioranoMQ, Hibernate, Swing, JMS, Maven, Ant, Oracle, Sybase, MS SQL Server, PostgreSQL, Caplin, Jabber, Microsoft Access, Teleforms,

• System Administration: GNU/Linux (Debian/Redhat), Novell Netware 3.12, Windows 9x, Windows 2000 I am an active user of Open Source software and enjoy the flexibility that it provides. I follow developments in the industry closely, and read several developers’ mailing-lists.

Certification

• Sun Certfied Programmer for the Java 1.4 platform

• Oracle 9i SQL Certification

Education October 2001 – June 2004 Magdalene College, University of Cambridge • Mathematics BA (Hons): First

September 1994 – June 2001 Cockermouth School, Cockermouth, Cumbria

• ‘A’-Levels: Mathematics (A), Further Mathematics (A), Physics (A), German (B), General Studies (A)

• STEP: Mathematics II (Grade S), Mathematics III (Grade 3)

• GCSEs: 7 A*s (including Mathematics, English Language, English Literature), 2 As

Other Interests

I have a passion for music—both listening to and playing it (I play the piano and the accordion to a good standard). I have been keen skier from a young age, enjoy travelling and amateur photography.

References Available on request.

This Curriculum Vitae was last updated on March 16, 2010.